OpenStudy (anonymous):

hi can somebody tell me what the formula is to work out a unlike fraction for exp 17/20 - 3/5 = 1/4

OpenStudy (anonymous):

take lcd of 20 and 5 = 20 (17(1) - 3(4))/20 =1/4

OpenStudy (anonymous):

thank you but could you make that a little easier because it was a little confusing sorry (:

OpenStudy (anonymous):

take lcd of 20 and 5 = the first number which 5 and 20 both have = 20 {17(number at which lcd is available at 20) - 3(number at which lcd is available at 5)}/20 {17(1) - 3(4)}/20

OpenStudy (anonymous):

You got it? @finmath

OpenStudy (anonymous):

yes thanks you @sangya21

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

There is one more 9/10 -1/15 = something something =

OpenStudy (anonymous):

check lcd of 10 and 15 = 30 (both have 30 in their sequence) 9(3)-1(2)/30 = 18-2/30